If f is continuous on [a,b] and f' is bounded in(a,b) prove that f is of bounded variation on [a,b]


1
Expert's answer
2021-03-19T04:47:55-0400

Since "f'" is bounded,

"\exist \, M>0 \ni | f'(\alpha)| \leq M \, \forall \alpha \in [a,b]"

Let "P = \{ x_0,x_1,...,x_n \}" be a partition of "[a,b]"

Then by the Mean Value Theorem, we choose "\alpha \in (x_{i-1},x_i) \, \ni" "f(x_i) - f(x_{i-1}) = f'(\alpha)(x_i - x_{i-1})"

Therefore"|f(x_i) - f(x_{i-1})| = |f'(\alpha)||x_i - x_{i-1}| \leq M|x_i - x_{i-1}|"

Hence, we have

"\sum_{i=1}^{n}|f(x_i) - f(x_{i-1})| \leq M \sum_{i=1}^{n} |x_i - x_{i-1}| = M(b-a)"

"V(f,a,b,P) \leq M(b-a)"

Since "M(b-a) > 0" we have that f is of bounded variation.